What is the empirical formula of an unknown organic gas molecule composed of carbon, hydrogen, and oxygen which undergoes combustion to produce 4.4g of CO2 and 2.7g H2O


1
Expert's answer
2021-04-19T04:35:38-0400

Ans :moles of carbon=4.4/44=0.1

Moles of hydrogen=2.7/18=0.15

Ratios

Carbon=0.1/0.1=1

Hydrogen=0.15/0.1=1.5

Empirical formula CxHy=CH2